OpenStudy (anonymous):

A series is given by the formula Tn=3+2n-n^2 (a) Find the 25th term. (b) Find n if the nth term is −252.

OpenStudy (anonymous):

a)\[T _{25}=3+2(25)-(25)^2\] Solve for T. b)\[-252=3+2n-n^2\] Move everything to the RHS(Right Hand Side) \[n^2-2n-255\] Solve the quadratic

OpenStudy (anonymous):

\[n^2−2n−255=0\]

OpenStudy (anonymous):

how to solve the quadratic

OpenStudy (anonymous):

the answer for b is 17 th term

OpenStudy (anonymous):

i dont know how to solve n^2 -2n

OpenStudy (anonymous):

Use the quadratic formula...

OpenStudy (anonymous):

\[n=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

Remember the nth term must be positive. You can't have a negative nth term.